OpenStudy (anonymous):

Please Help & Explain: Which set of numbers represents a Pythagorean triple? A. 6, 9, 12 B. 7, 10, 12 C. 16, 18, 25 D. 27, 36, 45

OpenStudy (solomonzelman):

Check each one of them. \(\large\color{black}{ 6^2+9^2=12^2 }\) \(\large\color{black}{ 7^2+10^2=12^2 }\) \(\large\color{black}{ 16^2+18^2=25^2 }\) \(\large\color{black}{ 27^2+36^2=45^2 }\)

OpenStudy (solomonzelman):

the shortest thing would be though, is to divide each of the sides in the 4th set, by 9. And if you know that \(\large\color{black}{ 3,4,5 }\) is a pythagorean triple, you are good to go.